OpenStudy (anonymous):

Find the value of x that makes the statement true. 43=16+3x

OpenStudy (anonymous):

(43-16)/3

OpenStudy (saifoo.khan):

3x = 43 - 16 3x = 27 x = 27/9 x = 3

OpenStudy (anonymous):

\[43=16+3x\]\[3x=27\]\[x=9\]